OpenStudy (anonymous):

4y=-x-5 -y=-4x-4 perpendicular parallel neither

OpenStudy (texaschic101):

-y = -4x - 4 --- multiply by -1 to make y positive y = 4x + 4 slope = 4 and y intercept = 4 4y = -x - 5 y = -1/4x - 5/4 slope = -1/4 and y intercept = -5/4 Perpendicular lines will have negative reciprocal slopes. 4 and -1/4 are negative reciprocal slopes....these lines are perpendicular